OpenStudy (anonymous):

what would the answer for this be and i need exact steps on how to do this please. x - + 1 = 0 3

OpenStudy (anonymous):

Multiply all terms on LHS and RHS by 3 x 3*---- + 3*1 = 3*0 3 this gives x + 3 = 0 subtracting 3 from both sides x + 3 - 3 = 0 - 3 x = - 3

OpenStudy (anonymous):

thank you.